OpenStudy (anonymous):

how to solve 7x+6 ≡ 3 mod 41 ?

OpenStudy (amistre64):

mod 41 means that we count on a circle or repetition from 1 to 41

OpenStudy (amistre64):

1 = 42 2 = 43 3 = 44 4 = 45 etc

OpenStudy (anonymous):

what is x than?

OpenStudy (amistre64):

0 = 41 1 = 42 2 = 43 3 = 44 4 = 45 . . . . . . 41 =82

OpenStudy (amistre64):

depends, what doe 3 mod41 equal?

OpenStudy (anonymous):

it's 3 however, according to the definition of congruence, both 3 and 7x+1 should have the same remainder

OpenStudy (amistre64):

good, but i think you meant 7x+6 right? and 3 mod 41 = 44 7x + 6 = 44 7x = 38 x = 38/7 but im sure that you want an integer value

OpenStudy (anonymous):

the answer is x=23, check with wolframaplha

OpenStudy (amistre64):

23 sounds reasonable yes :)

OpenStudy (anonymous):

but how do I solve it, show me a suggestion

OpenStudy (amistre64):

7(23) + 6 140+21+6 167 divided by 41 = 4 with a remainder of 3 which is how it looks when its is already solved

OpenStudy (amistre64):

3 mod 41 means that the number is divisible by 41 and has a remainder of 3

OpenStudy (anonymous):

yeah..I know, but show me how to solve it with Euclides

OpenStudy (amistre64):

with euclides? i got no idea what that method would be

OpenStudy (anonymous):

ok, but thanks anyways

OpenStudy (amistre64):

good luck :)